Solve for k: 5{(3-2k)+k}-6 = 2k-4(-3k-2)-12. Express your answer as a simplified fraction. PLEASE HELP!
Answer by fractalier(6550)

Okay, from
5{(3-2k)+k}-6 = 2k-4(-3k-2)-12
5(3 - 2k) + 5k - 6 = 2k + 12k + 8 - 12
15 - 10k + 5k - 6 = 2k + 12k + 8 - 12
9 - 5k = 14k - 4
13 = 19k
k = 13/19